Give the length and width of another rectangular field that has the same perimeter but a larger area.A rectangular field is 125 meters long and 75 meters wide.

Answer by dkppathak(439)

Give the length and width of another rectangular field that has the same perimeter but a larger area.A rectangular field is 125 meters long and 75 meters wide.
perimeter of rectangle =2(l+b)
=2(125+75)
=400 meters
area of rectangle =lxb= 125x75=9375 m^2
let another rectangle length is 110 and breath is 90 meters
perimeter will =2(l +b)
= 2(110+90)= 400 meters
arean of rectangle= lxb =110x90 = 9900 m^2
like so many example can be arranged for the mathematical application